HUBBLE SPACE TELESCOPE DAILY REPORT #1181 PERIOD COVERED: 0700 EST Daily Status Report as of 196/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5568 (WF/PC-2 Cycle 4 Calibration: Decon) The WF/PC-2 was used to make internal flat, K-spot, and dark calibration observations as part of a decontamination operation. Eleven observations were received and displayed. One observation was reported to show some peculiar bright steaks. The other observations were reported to look okay. 1.2 Completed WF/PC-2 5662 (Cycle 4 DD HST Observations of Comet Shoemaker-Levy (1993E) For July 1) The WF/PC-2 was used to make images of the comet Shoemaker-Levy (SL-5) to help refine the size estimates of the various components of the comet train. The observations were executed as scheduled. Eleven observations were received and displayed. They all were reported to look nominal. 1.3 Completed FOS/RD 5642 (Observations of the Impact of Comet Shoemaker-Levy(1993E) and Jupiter) The Faint Object Spectrograph (Red Side) made observations to support the Jupiter/Shoemaker-Levy event. Two observations were received and displayed in the OSS. They all were reported to look okay. 1.4 Completed Two Sets of WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1: Flats Part 2) The WF/PC-2 was used to make visual flats calibration images. Four images were received and displayed. One image was underexposed, one was partially saturated, and two were reported to be okay. 1.5 Completed FOS/RD 5455 (The Level of Ionization and Chemical Composition of QSO BL Region Gas) The Faint Object Spectrograph (Red Side) was used to make ultraviolet observations of quasi-stellar object (QSO) H1413+117 which is known to have broad absorption lines (BALs). The observations were executed as scheduled, and the data remains on Science Tape Recorder #1. It will be analyzed later. 1.6 Complete WF/PC-2 5476 (Continuum Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C298.0 as part of a continuum survey and to study the galaxy's morphology. The observations were executed as scheduled, and the data remains on Science Tape Recorder #1. It will be analyzed later. 1.7 Completed GHRS/2 5661 (FUV Spectroscopy of Io's SO2 Atmosphere and Surface Cycle 4 Carry-over) The High Resolution Spectrograph (Side 2) was used to make far ultraviolet spectrographic observations of Jupiter's moon Io (Shadow/Egress). The first onboard target acquisition was reported to be successful. The data remains on the science tape recorder, and will be analyzed later. 1.8 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of NGC-5597. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed FOS/RD 5456 (The BAL Region Covering Factor in QSOs with Weak Narrow-Line (O III) Emissions) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of the active galactic nuclei of nearby QSO Q1317-0142 to study weak, narrow O III emissions and strong Fe II emissions in BAL regions. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 5369 (HST Medium-Deep survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled :12 Successful :12 2.3 Significant HSTARs: None 2.4 Operations Notes: As the HST approaches the period of heaviest involvement in Comet Shoemaker-Levy activities, an anomaly-free operational day was experienced. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-